3) XO, Inc. reports the following cost information for March ost of Goods Manufactured anufacturing Overhead inished Goods Inventory, March 1 inished Goods Inventory, March 31 ork-in-Process Inventory, March 1 Vork-in-Process Inventorv, March 31 irect Materials Used 74,000 18,300 5000 880 1220 25,950 What is the amount of direct labor incurred by XO in March? C) $7650 A) $74,000 B) $21,090 D) $65,340
Answer #1

Cost of goods manufactured=Total manufacturing cost+Beginning WIP-Ending WIP

74000=Total manufacturing cost+9880-1220

Total manufacturing cost=(74000+1220-9880)=$65340

Total manufacturing cost=Direct labor+Direct material+Manufacturing overhead

65340=Direct labor+25950+18300

Direct labor=(65340-25950-18300)

which is equal to

=$21090.